# Discrete Diffusion VLA: Bringing Discrete Diffusion to Action Decoding in Vision-Language-Action Policies

### v2

@@ Abstract
Vision–Language–Action (VLA) models adapt large vision–language backbones to map images and instructions into robot actions. However, prevailing VLAs either generate actions autoregressively in a fixed left-to-right order or attach separate MLP or diffusion heads outside the backbone, leading to fragmented information pathways and specialized training requirements that hinder a unified, scalable architecture. We present Discrete Diffusion VLA, a unified-transformer policy that models discretized action chunks with discrete diffusion. The design retains diffusion’s progressive refinement paradigm while remaining natively compatible with the discrete token interface of VLMs. Our method achieves an adaptive decoding order that resolves easy action elements before harder ones and uses secondary re-masking to revisit uncertain predictions across refinement rounds, which improves consistency and enables robust error correction. This unified decoder preserves pretrained vision-language priors, supports parallel decoding, breaks the autoregressive bottleneck, and reduces the number of function evaluations. Discrete Diffusion VLA achieves 96.3% avg. success rates on LIBERO, 71.2% visual matching on SimplerEnv-Fractal and 54.2% overall on SimplerEnv–Bridge, improving over autoregressive, MLP decoder and continuous diffusion baselines. These findings indicate that discrete-diffusion VLA supports precise action modeling and consistent training, laying groundwork for scaling VLA to larger models and datasets. Our code will be available at https://github.com/Liang-ZX/DiscreteDiffusionVLA.
@@ 2.1 Vision-Language-Action Models
Early VLA systems took a two-part form: RT-1 and RT-2 first let the VLM produce latent tokens, then a separate MLP action decoder maps them to discretized controls in a single shot (Brohan et al., 2022; Zitkovich et al., 2023). Subsequent work shifted to AR token policies and scaled backbones and vision encoders for general manipulation (e.g. OpenVLA) (Kim et al., 2024; Touvron et al., 2023b; Oquab et al., 2024a; Zhai et al., 2023), while learning latent action shows better versatility (e.g. LAPA (Ye et al., 2024)). For high-frequency control and fast adaptation, token compression and action chunking are introduced and improve efficiency (e.g. $\pi_{0}$-FAST and OpenVLA-OFT) (Pertsch et al., 2025; Kim et al., 2025b; Song et al., 2025). In parallel, diffusion and flow matching action heads model continuous trajectories (Janner et al., 2022; Chi et al., 2023; Liang et al., 2023; Liu et al., 2024; Ni et al., 2023; Liang et al., 2025; Shen et al., 2025), spanning lightweight transformers (Li et al., 2024a), and hierarchical designs (e.g. $\pi_{0}$$\pi_{0.5}$(Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025b; a; Wen et al., 2025; Zhong et al., 2025; Liu et al., 2025; Liang et al., 2024; Yang et al., 2025b), typically via a separate denoising loop conditioned on language and vision tokens.
@@ 2.2 Discrete Diffusion Models
Discrete diffusion has recently achieved strong results on discrete data, notably tokenized images and natural language. Foundational work like D3PM (Austin et al., 2021) formalizes discrete diffusion as a Markov chain. Different from continuous diffusion models, it factorizes across positions, so each token (represented as a one-hot vector) is independently corrupted into a categorical distribution. VQ-Diffusion (Gu et al., 2022) and MaskGIT (Chang et al., 2022a) build on this, achieving high-fidelity image generation with transformers that iteratively predict masked/corrupted image tokens in a non-autoregressive manner. In language, Diffusion-BERT (He et al., 2022) and subsequent Masked Diffusion Models (Shi et al., 2024c; Zheng et al., 2024) demonstrate the viability of this approach and more recent LLaDA (Nie et al., 2025b) and DiffuLLaMA (Gong et al., 2024) scale it to 7B LMs competitive with AR baselines. MMaDA (Yang et al., 2025a) further develops this direction and shows the unified discrete diffusion model can jointly generate text and images. Concurrently, VPDD (He et al., 2024) explores discrete diffusion for policy learning via large-scale video-prediction pretraining and action fine-tuning, but it does not instantiate a unified VLA.
Our work extends this line to the VLA model: we perform discrete diffusion over tokenized action chunks, preserving language capabilities and VLM synergy, yielding competitive state-of-the-art VLA performance and laying groundwork toward unified, scalable, multi-modal foundation models spanning vision, language, and actions.
@@ 3.2 Discrete Diffusion over Action Tokens
$p_{\theta}(a_{t-1,i}\mid a_{t,i},\mathbf{c})=\begin{cases}\delta\big(a_{t-1,i}=a_{t,i}\big),&a_{t,i}\neq\mathrm{M},\\[4.0pt] \mathrm{Categorical}\big(a_{t-1,i}\mid\pi_{\theta}(i\mid\mathbf{c})\big),&a_{t,i}=\mathrm{M},\end{cases}$ (4)
@@
$\mathcal{L}_{\text{CE}}(\theta)\;=\;-\sum_{i\in\mathcal{M}_{\gamma_{t}}}\log p_{\theta}\bigl(a_{0,i}\mid\tilde{\mathbf{a}}_{t},\mathbf{c}\bigr),\qquad p_{\theta}(\cdot)=\mathrm{softmax}\bigl(W\,f_{\theta}(\tilde{\mathbf{a}}_{t},\mathbf{c})\bigr),$ (5)
@@ 3.3 Our Unified VLA Architecture
Compared to prior two action head designs, this unified backbone retains the language model’s representation ability, scales seamlessly with model size, and allows parallel decoding. At inference time, our adaptive re-masking (detailed in Sec. 3.5) further refines uncertain tokens, combining the global-context strength of transformers with the iterative denoising spirit of diffusion.
@@ 4.1 Benchmarks and Baselines
OpenVLA-OFT (Cont-Diffusion) 96.9 98.1 95.5 91.1 95.4
@@ 4.3 Ablation Study
Decoding Strategy Parallel Decoding Random Order Confidence Gap Max Confidence Max Conf+ Second Remask
Success 95.6% 96.0% 96.6% 97.0% 97.4%
@@ 4.4 Analysis on Inference Efficiency
Speed–quality trade-off. Figure 4 sweeps $T$ and reports throughput (number of action chunks generated per second; left $y$-axis) and success rates on LIBERO-GOAL (right $y$-axis). Accuracy improves monotonically with $T$ but with diminishing returns: most gains appear by $T{\in}[8,12]$, while increasing to $T{\geq}16$ yields sub–1 point improvements at a disproportionate drop in speed (throughput scales roughly as $1/T$). We therefore choose T=12T{=}12 as a knee point that matches the best accuracy in Tables 45 while maintaining high throughput (about 3 Hz for action chunk generation).
@@ Acknowledgments
We sincerely thank Wanqi Zhong from Harbin Institute of Technology for assistance with figure illustration and layout. This project is partially supported by Chinese Institute of Electronics-Tencent Doctoral Research Incentive Program.

### v3

@@ Abstract
Vision–Language–Action (VLA) models adapt large vision–language backbones to map images and instructions into robot actions. However, prevailing VLAs either generate actions autoregressively in a fixed left-to-right order or attach separate MLP or diffusion heads outside the backbone, leading to fragmented information pathways and specialized training requirements that hinder a unified, scalable architecture. We present Discrete Diffusion VLA, a unified-transformer policy that models discretized action chunks with discrete diffusion. The design retains diffusion’s progressive refinement paradigm while remaining natively compatible with the discrete token interface of VLMs. Our method achieves an adaptive decoding order that resolves easy action elements before harder ones and uses secondary re-masking to revisit uncertain predictions across refinement rounds, which improves consistency and enables robust error correction. This unified decoder preserves pretrained vision-language priors, supports parallel decoding, breaks the autoregressive bottleneck, and reduces the number of function evaluations. Discrete Diffusion VLA achieves 96.3% avg. success rates on LIBERO, 71.2% visual matching on SimplerEnv-Fractal and 54.2% overall on SimplerEnv–Bridge. We also provide ablation study on vision-language ability retention on LIBERO-OOD (Out-of-Distribution) benchmark, with our method improving over autoregressive, MLP decoder and continuous diffusion baselines. These findings indicate that discrete-diffusion VLA supports precise action modeling and consistent training, laying groundwork for scaling VLA to larger models and datasets. Our code is available at this page.
@@ 2.1 Vision-Language-Action Models
Early VLA systems took a two-part form: RT-1 and RT-2 first let the VLM produce latent tokens, then a separate MLP action decoder maps them to discretized controls in a single shot (Brohan et al., 2022; Zitkovich et al., 2023). Subsequent work shifted to AR token policies and scaled backbones and vision encoders for general manipulation (e.g. OpenVLA) (Kim et al., 2024; Touvron et al., 2023b; Oquab et al., 2024a; Zhai et al., 2023), while learning latent action shows better versatility (e.g. LAPA (Ye et al., 2024)). For high-frequency control and fast adaptation, token compression and action chunking are introduced and improve efficiency (e.g. $\pi_{0}$-FAST and OpenVLA-OFT) (Pertsch et al., 2025; Kim et al., 2025b). In parallel, diffusion and flow matching action heads model continuous trajectories (Janner et al., 2022; Chi et al., 2023; Liang et al., 2023; Liu et al., 2024), spanning lightweight transformers (Li et al., 2024a), and hierarchical designs (e.g. $\pi_{0}$$\pi_{0.5}$(Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025b; a; Wen et al., 2025; Zhong et al., 2025; Liu et al., 2025), typically via a separate denoising loop conditioned on language and vision tokens.
@@ 2.2 Discrete Diffusion Models
Discrete diffusion has recently achieved strong results on discrete data, notably tokenized images and natural language. Foundational work like D3PM (Austin et al., 2021) formalizes discrete diffusion as a Markov chain. Different from continuous diffusion models, it factorizes across positions, so each token (represented as a one-hot vector) is independently corrupted into a categorical distribution. VQ-Diffusion (Gu et al., 2022) and MaskGIT (Chang et al., 2022a) build on this, achieving high-fidelity image generation with transformers that iteratively predict masked/corrupted image tokens in a non-autoregressive manner. In language, Diffusion-BERT (He et al., 2022) and subsequent Masked Diffusion Models (Shi et al., 2024c; Zheng et al., 2024) demonstrate the viability of this approach and more recent LLaDA (Nie et al., 2025b) and DiffuLLaMA (Gong et al., 2024) scale it to 7B LMs competitive with AR baselines. MMaDA (Yang et al., 2025a) further develops this direction and shows the unified discrete diffusion model can jointly generate text and images.
Our work extends this line to the action modality: we perform discrete diffusion over tokenized action chunks, preserving language capabilities and VLM synergy, yielding competitive state-of-the-art VLA performance and laying groundwork toward unified, scalable, multi-modal foundation models spanning vision, language, and actions.
@@ 3.2 Discrete Diffusion over Action Tokens
$\vskip-2.0ptp_{\theta}(a_{t-1,i}\mid a_{t,i},\mathbf{c})=\begin{cases}\delta\big(a_{t-1,i}=a_{t,i}\big),&a_{t,i}\neq\mathrm{M},\\[4.0pt] \mathrm{Categorical}\big(a_{t-1,i}\mid\pi_{\theta}(i\mid\mathbf{c})\big),&a_{t,i}=\mathrm{M},\end{cases}$ (4)
@@
$\vskip-3.0pt\mathcal{L}_{\text{CE}}(\theta)\;=\;-\sum_{i\in\mathcal{M}_{\gamma_{t}}}\log p_{\theta}\bigl(a_{0,i}\mid\tilde{\mathbf{a}}_{t},\mathbf{c}\bigr),\qquad p_{\theta}(\cdot)=\mathrm{softmax}\bigl(W\,f_{\theta}(\tilde{\mathbf{a}}_{t},\mathbf{c})\bigr),\vskip-2.0pt$ (5)
@@ 3.3 Our Unified VLA Architecture
Compared to prior two action head designs, this unified backbone retains the language model’s representation power, scales seamlessly with model size, and allows parallel decoding. At inference time, our adaptive re-masking (detailed in Sec. 3.5) further refines uncertain tokens, combining the global-context strength of transformers with the iterative denoising spirit of diffusion.
@@ 4.2 Overall Performance Comparisons
OpenVLA-OFT (Cont-Diffusion) 96.9 98.1 96.2 91.1 95.6
@@
4.3 Quantitative Analysis on Retention of Pretrained Vision-Language Priors
[…]
Figure 4: Vision Augmentation Sample of LIBERO-OOD Goal Task. The out-of-distribution (OOD) setting introduces visual perturbations by replacing objects with variants that differ in scale, material properties, and visual appearance. Specifically, the modified scenes include a larger bowl and a stove with metallic luster reflections.
Results. Tables 4 and 5 present results on LIBERO-Goal and LIBERO-Spatial, respectively.
Table 4: Out-of-distribution performance on LIBERO-Goal
Method Original Lang Aug Vision Aug
OpenVLA-OFT (Discrete) 95.6% 87.6% ($\downarrow$8.0%) 73.0% ($\downarrow$22.6%)
OpenVLA-OFT (Cont-Diffusion) 96.0% 93.6% ($\downarrow$2.4%) 67.0% ($\downarrow$29.0%)
Discrete Diffusion VLA 97.4% 96.0% ($\downarrow$1.4%) 76.4% ($\downarrow$21.0%)
Table 5: Out-of-distribution performance on LIBERO-Spatial
Method Original Lang Aug Vision Aug
OpenVLA-OFT (Discrete) 96.2% 94.6% ($\downarrow$1.6%) 95.0% ($\downarrow$1.2%)
OpenVLA-OFT (Cont-Diffusion) 96.9% 95.2% ($\downarrow$1.7%) 94.6% ($\downarrow$2.3%)
Discrete Diffusion VLA 97.2% 96.0% ($\downarrow$1.2%) 96.2% ($\downarrow$1.0%)
Our method exhibits substantially smaller degradation under both language and vision perturbations. On LIBERO-Goal, language degradation is only 1.4% vs. 8.0% (parallel decoding) and 2.4% (continuous diffusion), while vision degradation is 21.0% vs. 22.6% and 29.0%. On LIBERO-Spatial, degradations are minimal across all methods ($\leq$3.0%), reflecting the suite’s inherent layout variations during training. The stark contrast with LIBERO-Goal stems from its limited task diversity—ten sparse tasks with large goal-space gaps that encourage overfitting to specific instances rather than generalizable understanding.
Architectural Insights. Parallel decoding shows higher language sensitivity, likely due to PaliGemma’s language-centric modality alignment. Continuous diffusion exhibits severe vision degradation ($\downarrow$29.0% on LIBERO-Goal), consistent with prior findings (Yang et al., 2025b; Liu et al., 2024) that separate diffusion action heads become overly vision-dependent. Our unified discrete diffusion framework mitigates both issues through tighter VLM integration, demonstrating superior preservation of pretrained vision-language capabilities.
@@ 4.4 Ablation Study
Decoding Strategy Parallel Decoding +Random Order +Confidence Gap +Max Confidence +Max Conf. +Second Remask
Success 95.6% 95.8% 96.6% 97.0% 97.4%
@@
4.5 Visualization and Analysis of Adaptive Decoding Order
To validate Discrete Diffusion VLA’s “easy-first, hard-later” pattern, we visualized the decoding process across 12 denoising steps (Figure 6(a)). We use token frequency in the training data as a proxy for action difficulty, i.e.high-frequency tokens are generally easier to learn. In the visualization, light gray indicates masked tokens, while brown/orange shades represent demasked tokens, with darker colors indicating higher frequency.
Figure 6: Visualization and Analysis of Adaptive Decoding Order
The visualization reveals that darker colors (high-frequency tokens) predominantly appear in early timesteps ($t=0\sim t=6$), supporting our intuition that easier elements are resolved first. However, the pattern still follows trajectory structure rather than blindly prioritizing all high-frequency tokens, indicating adaptation to trajectory-specific confidence rather than dataset bias.
Figure 6(c) quantifies this at $t=4$: among 8 decoded tokens, 6 come from the top-5 most frequent tokens in the dataset (Table 6(b)), providing empirical evidence that our adaptive strategy prioritizes easy action elements during early refinement. More detailed analyses are show in Appendix E.
@@ 4.6 Analysis on Inference Efficiency
[…]
Table 8: Inference speed comparison on LIBERO-Goal.
Method Latency (ms) Speed (Hz) NFE
OpenVLA (AR) 136.2 7.34 56
OpenVLA w/o KVcache (AR) 209.5 4.77 56
OpenVLA-OFT (Parallel Decoding) 31.1 32.14 1
OpenVLA-OFT (Cont-Diffusion, 50 steps) 199.9 5.00 50
OpenVLA-OFT (Cont-Diffusion+, 12 steps) 67.1 14.91 12
Discrete Diffusion VLA (12 steps) 68.8 14.53 12
Speed–quality trade-off. Figure 5 sweeps $T$ and reports throughput (number of action chunks generated per second; left $y$-axis) and success rates on LIBERO-GOAL (right $y$-axis). Accuracy improves monotonically with $T$ but with diminishing returns: most gains appear by $T{\in}[8,12]$, while increasing to $T{\geq}16$ yields sub–1 point improvements at a disproportionate drop in speed (throughput scales roughly as $1/T$). We choose T=12T{=}12 as the optimal operating point balancing accuracy and efficiency.
@@ Acknowledgments
We sincerely thank Wanqi Zhong from Harbin Institute of Technology for assistance with figure illustration and layout.
@@ Appendix C Implementation Details
Appendix D Out-of Distribution Evaluation Benchmark
We follow the general experimental settings of LIBERO-PRO (Zhou et al., 2025) while correcting several inconsistencies identified in the currently published implementation.
D.1 LIBERO-GOAL-OOD
The out-of-distribution (OOD) setting introduces visual perturbations by replacing objects with variants that differ in scale, material properties, and visual appearance. Specifically, the modified scenes include a larger bowl and a stove with metallic luster reflections.
Another case with a bottle and a cabinet constructed from different materials and textures to further challenge the model’s generalization capabilities.
Examples of augmented instructions:
  • slice open the top drawer and place the bowl in it
  • open the middle dresser of the cabinet
  • place the bowl on top of the cabinet
D.2 LIBERO-SPATIAL-OOD
The visual out-of-distribution setting is similar to that of LIBERO-GOAL-OOD with different scale, material property, and visual appearance objects.
Examples of augmented instructions:
  • lift the black bowl on the stove and set it on the plate
  • take the black bowl beside the plate and put it on the plate
  • grab the black bowl on the ramekin and set it on the plate
Appendix E Visualization and Analysis of Adaptive Decoding Order
To validate the intuition of Discrete Diffusion VLA’s “easy-first, hard-later” decoding pattern, we conducted a comprehensive analysis of the learned adaptive decoding order. We visualized the decoding process for an action chunk across 12 denoising steps, as shown in Figure 6 (a). Each grid represents an $8\times 7$ layout, where rows correspond to 8 actions in a chunk and columns represent 7-dimensional tokens per action.
We additionally computed the frequency distribution of action tokens in the fine-tuning dataset. Table 6 (b) shows the top 5 most frequent token IDs for each action dimension, with their occurrence percentages. Then in Fig. 6 (a), light gray indicates masked tokens, while varying shades of brown / orange represent demasked tokens, with color intensity encoding token frequency in the training dataset. (Darker colors mean high frequency.)
From the figure, we can find that darker colors (high-frequency tokens) are predominantly decoded in the early timesteps ($t=0\sim t=6$). While token frequency does not directly equal confidence, it can serve as a reasonable proxy for easy / hard action elements because tokens that appear more frequently in training tend to be learned more robustly by the model. This generally supports our claimed intuition that the adaptive decoding mechanism tends to resolve easier action elements before harder ones, though the correspondence is not absolute.
Besides, while high-frequency tokens show a tendency to be decoded earlier, the pattern still follows the actual trajectory structure that not all dark tokens are decoded first. This indicates our strategy adapts to trajectory-specific confidence rather than blindly following dataset statistics, avoiding collapse to dataset bias.
Figure 6 (c) further shows the specific decoded tokens at $t=4$. Among the 8 decoded tokens, 6 out of 8 come from the top-5 most frequent tokens in Table 6 (b), providing quantitative evidence that our adaptive strategy often prioritizes easy action elements in early refinement steps.